OpenStudy (marmar10):

Covert 1.53×103 kg/ML to Mg/cL

OpenStudy (kittiwitti1):

Oooh conversions! :o Is 103 actually 10^3?

OpenStudy (marmar10):

yes oops

OpenStudy (kittiwitti1):

Okay. I can't do the work for you because then you wouldn't learn, but here is an acronym for you to help remember unit conversions: \[\text{Great Mighty king henry Died __ drinking chocolate milk monday night}\]\[\text{Giga Mega kilo hecto Deca (base unit) deci centi milli micro nano}\]\[G-M-k-h-D-\text{base unit}-d-c-m-\mu-n\]

OpenStudy (kittiwitti1):

https://www.mnemonic-device.com/physics/metric/prefixes-of-the-metric-system/ This link can help you.

OpenStudy (kittiwitti1):

Conversion formula (input values based on the link/template I gave you before)\[\frac{1.53\times10^{3}kg}{ML}=\frac{?g}{1kg}=\frac{1Mg}{?g}=\frac{?L}{1ML}=\frac{?cL}{1L}\]

OpenStudy (marmar10):

thank you! i will try to do it now

OpenStudy (kittiwitti1):

Alright then :] You're welcome ^^ feel free to come back here if you run into any trouble doing it, I'll be here :p

OpenStudy (marmar10):

i got .0000153 but its not right

OpenStudy (kittiwitti1):

Did you put it in scientific notation?

OpenStudy (marmar10):

still not right :/

OpenStudy (marmar10):

is that even the right answer?

OpenStudy (kittiwitti1):

Hmm. Let me try.

OpenStudy (kittiwitti1):

You're not on a time limit are you? I need awhile to work this out lol

OpenStudy (marmar10):

nope!

OpenStudy (kittiwitti1):

Alright good. Brb then xD

OpenStudy (kittiwitti1):

._. Oh heck... I messed up the formula. That's probably why you got it wrong lol

OpenStudy (kittiwitti1):

\[\frac{1.53\times10^{3}kg}{1ML}=\frac{?g}{1kg}=\frac{1Mg}{?g}=\frac{1ML}{?L}=\frac{?cL}{1L}\]Can you do it now?

OpenStudy (marmar10):

hmmm still wrong

OpenStudy (kittiwitti1):

What answer did you get?

OpenStudy (marmar10):

15.3

OpenStudy (kittiwitti1):

Did your equation look like this?\[\frac{1.53\times10^{3}kg}{1ML}=\frac{10^{3}g}{1kg}=\frac{1ML}{10^{6}L}=\frac{10^{2}cL}{1L}\]

OpenStudy (marmar10):

nope. you forgot 1Mg/g

OpenStudy (marmar10):

i only have one more try lol :')

OpenStudy (kittiwitti1):

Oh right sorry! \[=\frac{1Mg}{10^{6}g}\]What about now?

OpenStudy (kittiwitti1):

Don't try again, do it on paper >_>

OpenStudy (kittiwitti1):

OH WOW I did the formula wrong again =_= flip the last fraction in this http://prntscr.com/8kv07d

OpenStudy (marmar10):

got .00153

OpenStudy (kittiwitti1):

Okay so you get \[\frac{1.53\times10^{3}}{1ML}=\frac{10^{3}g}{1kg}=\frac{1Mg}{10^{6}}=\frac{1ML}{10^{6}L}=\frac{1L}{10^{2}cL}\]

OpenStudy (kittiwitti1):

cross out like units then simplify:\[\frac{1.53\times10^{3}\times10^{3}Mg}{10^{6}\times10^{6}\times10^{2}cL}\]what did you get?

OpenStudy (marmar10):

.0000000153

OpenStudy (kittiwitti1):

it's 10^3 on top and 10^6, 10^6, 10^2 on bottom btw.\[\frac{1.53\times10^{6}}{10^{14}}\]

OpenStudy (kittiwitti1):

\[1.53\times10^{-8}\]Did you get this?

OpenStudy (marmar10):

yep same thing!

OpenStudy (kittiwitti1):

Alright then I think that's the right answer. You still have one attempt left right?

OpenStudy (marmar10):

yay its right! haha thank you. conversions suck

OpenStudy (kittiwitti1):

Lol, honestly it would've been easier if I'd written it out from the start xD I messed up the calculations in my head e_e